a body in a straight line travels 2 m in the first two seconds and 2.2m in the next four seconds with constant retardation. what will be its velocity at the end of the seventh second from the start?

Grade:9

2 Answers

Sunil Kumar FP
9 years ago
let the initial velocity be u
velocity at the end of second sth second=v1
velocity at the end of 6thsecond=v2
velocity at the end of the 7thsecond=v3
retardation=@
v1=u-2@
v2=u-6@
v3=u-7@
now
2=(u-2@)2-1/2@4
u-@=1
2.2=(u-2@)4-1/2@16
@=3/20
now v3=u-7@
=0.1m/s
thanks and regards
sunil kr
askIITian faculty
Sumit Majumdar IIT Delhi
9 years ago
Dear student,
From the given information, we have:
s_{2}=2u+2a=2 ...(i)
s_{6}-s_{2}=6u+18a-2u-2a=4u+16a=2.2 ...(ii)
Here s reperesents the displacement, u the initial velocity, a is the acceleration.
Hence, we get: a = -0.15 m/s2, u = 1.15 m/s.
Thus, we get:
s_{7}=7u+24.5a=4.375 m
Regards
Sumit